### v2

# Graph Retrieval-Augmented Generation: A Survey

Abstract.
Recently, Retrieval-Augmented Generation (RAG) has achieved remarkable success in addressing the challenges of Large Language Models (LLMs) without necessitating retraining. By referencing an external knowledge base, RAG refines LLM outputs, effectively mitigating issues such as “hallucination”, lack of domain-specific knowledge, and outdated information. However, the complex structure of relationships among different entities in databases presents challenges for RAG systems. In response, GraphRAG leverages structural information across entities to enable more precise and comprehensive retrieval, capturing relational knowledge and facilitating more accurate, context-aware responses. Given the novelty and potential of GraphRAG, a systematic review of current technologies is imperative. This paper provides the first comprehensive overview of GraphRAG methodologies. We formalize the GraphRAG workflow, encompassing Graph-Based Indexing, Graph-Guided Retrieval, and Graph-Enhanced Generation. We then outline the core technologies and training methods at each stage. Additionally, we examine downstream tasks, application domains, evaluation methodologies, and industrial use cases of GraphRAG. Finally, we explore future research directions to inspire further inquiries and advance progress in the field. In order to track recent progress in this field, we set up a repository at https://github.com/pengboci/GraphRAG-Survey.
Large Language Models, Graph Retrieval-Augmented Generation, Knowledge Graphs, Graph Neural Networks
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 9ccs: Computing methodologies Knowledge representation and reasoningccs: Information systems Information retrievalccs: Information systems Data mining
1. Introduction
The development of Large Language Models like GPT-4 (OpenAI, 2024), Qwen2 (Yang et al., 2024c), and LLaMA (Dubey et al., 2024) has sparked a revolution in the field of artificial intelligence, fundamentally altering the landscape of natural language processing. These models, built on Transformer (Vaswani et al., 2017) architectures and trained on diverse and extensive datasets, have demonstrated unprecedented capabilities in understanding, interpreting, and generating human language. The impact of these advancements is profound, stretching across various sectors including healthcare (Wang et al., 2024c; Liu et al., 2024c; Zheng et al., 2024), finance (Nie et al., 2024; Li et al., 2024d), and education (Wang et al., 2024d; Ghimire et al., 2024), where they facilitate more nuanced and efficient interactions between humans and machines.
Despite their remarkable language comprehension and text generation capabilities, LLMs may exhibit limitations due to a lack of domain-specific knowledge, real-time updated information, and proprietary knowledge, which are outside LLMs’ pre-training corpus. These gaps can lead to a phenomenon known as “hallucination” (Huang et al., 2023b) where the model generates inaccurate or even fabricated information. Consequently, it is imperative to supplement LLMs with external knowledge to mitigate this problem. Retrieval-Augmented Generation (RAG) (Fan et al., 2024a; Hu and Lu, 2024; Huang and Huang, 2024; Zhao et al., 2024; Gao et al., 2024b; Yu et al., 2024; Wu et al., 2024b) emerged as a significant evolution, which aims to enhance the quality and relevance of generated content by integrating a retrieval component within the generation process. The essence of RAG lies in its ability to dynamically query a large text corpus to incorporate relevant factual knowledge into the responses generated by the underlying language models. This integration not only enriches the contextual depth of the responses but also ensures a higher degree of factual accuracy and specificity. RAG has gained widespread attention due to its exceptional performance and broad applications, becoming a key focus within the field.
Figure 1. Comparision between Direct LLM, RAG, and GraphRAG. Given a user query, direct answering by LLMs may suffer from shallow responses or lack of specificity. RAG addresses this by retrieving relevant textual information, somewhat alleviating the issue. However, due to the text’s length and flexible natural language expressions of entity relationships, RAG struggles to emphasize “influence” relations, which is the core of the question. While, GraphRAG methods leverage explicit entity and relationship representations in graph data, enabling precise answers by retrieving relevant structured information.
Although RAG has achieved impressive results and has been widely applied across various domains, it faces limitations in real-world scenarios: (1) Neglecting Relationships: In practice, textual content is not isolated but interconnected. Traditional RAG fails to capture significant structured relational knowledge that cannot be represented through semantic similarity alone. For instance, in a citation network where papers are linked by citation relationships, traditional RAG methods focus on finding the relevant papers based on the query but overlook important citation relationships between papers. (2) Redundant Information: RAG often recounts content in the form of textual snippets when concatenated as prompts. This makes context become excessively lengthy, leading to the “lost in the middle” dilemma (Liu et al., 2024a). (3) Lacking Global Information: RAG can only retrieve a subset of documents and fails to grasp global information comprehensively, and hence struggles with tasks such as Query-Focused Summarization (QFS).
Graph Retrieval-Augmented Generation (GraphRAG) (Edge et al., 2024; Hu et al., 2024; Mavromatis and Karypis, 2024) emerges as an innovative solution to address these challenges. Unlike traditional RAG, GraphRAG retrieves graph elements containing relational knowledge pertinent to a given query from a pre-constructed graph database, as depicted in Figure 1. These elements may include nodes, triples, paths, or subgraphs, which are utilized to generate responses. GraphRAG considers the interconnections between texts, enabling a more accurate and comprehensive retrieval of relational information. Additionally, graph data, such as knowledge graphs, offer abstraction and summarization of textual data, thereby significantly shortening the length of the input text and mitigating concerns of verbosity. By retrieving subgraphs or graph communities, we can access comprehensive information to effectively address the QFS challenge by capturing the broader context and interconnections within the graph structure.
In this paper, we are the first to provide a systematic survey of GraphRAG. Specifically, we begin by introducing the GraphRAG workflow, along with the foundational background knowledge that underpins the field. Then, we categorize the literature according to the primary stages of the GraphRAG process: Graph-Based Indexing (G-Indexing), Graph-Guided Retrieval (G-Retrieval), and Graph-Enhanced Generation (G-Generation) in Section 5, Section 6 and Section 7 respectively, detailing the core technologies and training methods within each phase. Furthermore, we investigate downstream tasks, application domains, evaluation methodologies, and industrial use cases of GraphRAG. This exploration elucidates how GraphRAG is being utilized in practical settings and reflects its versatility and adaptability across various sectors. Finally, acknowledging that research in GraphRAG is still in its early stages, we delve into potential future research directions. This prognostic discussion aims to pave the way for forthcoming studies, inspire new lines of inquiry, and catalyze progress within the field, ultimately propelling GraphRAG toward more mature and innovative horizons.
Our contributions can be summarized as follows:
  • We provide a comprehensive and systematic review of existing state-of-the-art GraphRAG methodologies. We offer a formal definition of GraphRAG, outlining its universal workflow which includes G-Indexing, G-Retrieval, and G-Generation.
  • We discuss the core technologies underpinning existing GraphRAG systems, including G-Indexing, G-Retrieval, and G-Generation. For each component, we analyze the spectrum of model selection, methodological design, and enhancement strategies currently being explored. Additionally, we contrast the diverse training methodologies employed across these modules.
  • We delineate the downstream tasks, benchmarks, application domains, evaluation metrics, current challenges, and future research directions pertinent to GraphRAG, discussing both the progress and prospects of this field. Furthermore, we compile an inventory of existing industry GraphRAG systems, providing insights into the translation of academic research into real-world industry solutions.
Organization. The rest of the survey is organized as follows: Section 2 compares related techniques, while Section 3 outlines the general process of GraphRAG. Sections 5 to 7 categorize the techniques associated with GraphRAG’s three stages: G-Indexing, G-Retrieval, and G-Generation. Section 8 introduces the training strategies of retrievers and generators. Section 9 summarizes GraphRAG’s downstream tasks, corresponding benchmarks, application domains, evaluation metrics, and industrial GraphRAG systems. Section 10 provides an outlook on future directions. Finally, Section 11 concludes the content of this survey.
Figure 2. The overview of the GraphRAG framework for question answering task. In this survey, we divide GraphRAG into three stages: G-Indexing, G-Retrieval, and G-Generation. We categorize the retrieval sources into open-source knowledge graphs and self-constructed graph data. Various enhancing techniques like query enhancement and knowledge enhancement may be adopted to boost the relevance of the results. Unlike RAG, which uses retrieved text directly for generation, GraphRAG requires converting the retrieved graph information into patterns acceptable to generators to enhance the task performance.
2. Comparison with Related Techniques and Surveys
In this section, we compare Graph Retrieval-Augmented Generation (GraphRAG) with related techniques and corresponding surveys, including RAG, LLMs on graphs, and Knowledge Base Question Answering (KBQA).
2.1. RAG
RAG combines external knowledge with LLMs for improved task performance, integrating domain-specific information to ensure factuality and credibility. In the past two years, researchers have written many comprehensive surveys about RAG (Fan et al., 2024a; Hu and Lu, 2024; Huang and Huang, 2024; Zhao et al., 2024; Gao et al., 2024b; Yu et al., 2024; Wu et al., 2024b). For example,  Fan et al. (2024a) and Gao et al. (2024b) categorize RAG methods from the perspectives of retrieval, generation, and augmentation.  Zhao et al. (2024) review RAG methods for databases with different modalities.  Yu et al. (2024) systematically summarize the evaluation of RAG methods. These works provide a structured synthesis of current RAG methodologies, fostering a deeper understanding and suggesting future directions of the area.
From a broad perspective, GraphRAG can be seen as a branch of RAG, which retrieves relevant relational knowledge from graph databases instead of text corpus. However, compared to text-based RAG, GraphRAG takes into account the relationships between texts and incorporates the structural information as additional knowledge beyond text. Furthermore, during the construction of graph data, raw text data may undergo filtering and summarization processes, enhancing the refinement of information within the graph data. Although previous surveys on RAG have touched upon GraphRAG, they predominantly center on textual data integration. This paper diverges by placing a primary emphasis on the indexing, retrieval, and utilization of structured graph data, which represents a substantial departure from handling purely textual information and spurs the emergence of many new techniques.
2.2. LLMs on Graphs
LLMs are revolutionizing natural language processing due to their excellent text understanding, reasoning, and generation capabilities, along with their generalization and zero-shot transfer abilities. Although LLMs are primarily designed to process pure text and struggle with non-Euclidean data containing complex structural information, such as graphs (Wang et al., 2023b; Guo et al., 2023), numerous studies (Fan et al., 2024b; Mao et al., 2024c; Liu et al., 2024d; Li et al., 2024c; Pan et al., 2023, 2024; Jin et al., 2024a; Chen, 2024; Zhu et al., 2024; Wang et al., 2024e) have been conducted in these fields. These papers primarily integrate LLMs with GNNs to enhance modeling capabilities for graph data, thereby improving performance on downstream tasks such as node classification, edge prediction, graph classification, and others. For example,  Zhu et al. (2024) propose an efficient fine-tuning method named ENGINE, which combines LLMs and GNNs through a side structure for enhancing graph representation.
Different from these methods, GraphRAG focuses on retrieving relevant graph elements using queries from an external graph-structured database. In this paper, we provide a detailed introduction to the relevant technologies and applications of GraphRAG, which are not included in previous surveys of LLMs on Graphs.
2.3. KBQA
KBQA is a significant task in natural language processing, aiming to respond to user queries based on external knowledge bases (Fu et al., 2020; Lan et al., 2023, 2021; Yani and Krisnadhi, 2021), thereby achieving goals such as fact verification, passage retrieval enhancement, and text understanding. Previous surveys typically categorize existing KBQA approaches into two main types: Information Retrieval (IR)-based methods and Semantic Parsing (SP)-based methods. Specifically, IR-based methods (Luo et al., 2024b; Sun et al., 2024b; Ma et al., 2024; Jiang et al., 2023b; Zhang et al., 2022b; Wu et al., 2023b; Wang et al., 2023a; Jiang et al., 2024b) retrieve information related to the query from the knowledge graph (KG) and use it to enhance the generation process. While SP-based methods (Chakraborty, 2024; Fang et al., 2024b; Chen et al., 2021; Ye et al., 2021; Gu and Su, 2022; Sun et al., 2023) generate a logical form (LF) for each query and execute it against knowledge bases to obtain the answer.
GraphRAG and KBQA are closely related, with IR-based KBQA methods representing a subset of GraphRAG approaches focused on downstream applications. In this work, we extend the discussion beyond KBQA to include GraphRAG’s applications across various downstream tasks. Our survey provides a thorough and detailed exploration of GraphRAG technology, offering a comprehensive understanding of existing methods and potential improvements.
3. Preliminaries
In this section, we introduce background knowledge of GraphRAG for easier comprehension of our survey. First, we introduce Text-Attributed Graphs which is a universal and general format of graph data used in GraphRAG. Then, we provide formal definitions for two types of models that can be used in the retrieval and generation stages: Graph Neural Networks and Language Models.
3.1. Text-Attributed Graphs
The graph data used in Graph RAG can be represented uniformly as Text-Attributed Graphs (TAGs), where nodes and edges possess textual attributes. Formally, a text-attributed graph can be denoted as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{A},\{\mathbf{x}_{v}\}_{v\in\mathcal{V}},\{\mathbf{e}_{i,j}\}_{i,j\in\mathcal{E}})$, where $\mathcal{V}$ is the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges, $\mathcal{A}\in\{0,1\}^{\lvert\mathcal{V}\rvert\times\lvert\mathcal{V}\rvert}$ is the adjacent matrix. Additionally, $\{\mathbf{x}_{v}\}_{v\in\mathcal{V}}$ and $\{\mathbf{e}_{i,j}\}_{i,j\in\mathcal{E}}$ are textual attributes of nodes and edges, respectively. One typical kind of TAGs is Knowledge Graphs (KGs), where nodes are entities, edges are relations among entities, and text attributes are the names of entities and relations.
3.2. Graph Neural Networks
Graph Neural Networks (GNNs) are a kind of deep learning framework to model the graph data. Classical GNNs, e.g., GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), GraphSAGE (Hamilton et al., 2017), adopt a message-passing manner to obtain node representations. Formally, each node representation $\mathbf{h}^{(l-1)}_{i}$ in the $l$-th layer is updated by aggregating the information from neighboring nodes and edges:
(1) $\mathbf{h}^{(l)}_{i}=\textbf{UPD}(\mathbf{h}^{(l-1)}_{i},\textbf{AGG}_{j\in\mathcal{N}(i)}\textbf{MSG}(\mathbf{h}^{(l-1)}_{i},\mathbf{h}^{(l-1)}_{j},\mathbf{e}^{(l-1)}_{i,j})),$
where $\mathcal{N}_{(i)}$ represents the neighbors of node $i$. MSG denotes the message function, which computes the message based on the node, its neighbor, and the edge between them. AGG refers to the aggregation function that combines the received messages using a permutation-invariant method, such as mean, sum, or max. UPD represents the update function, which updates each node’s attributes with the aggregated messages.
Subsequently, a readout function, e.g., mean, sum, or max pooling, can be applied to obtain the global-level representation:
(2) $\mathbf{h}_{G}=\textbf{READOUT}_{i\in\mathcal{V}_{G}}(\mathbf{h}_{i}^{(L)}).$
In GraphRAG, GNNs can be utilized to obtain representations of graph data for the retrieval phase, as well as to model the retrieved graph structures.
3.3. Language Models
Language models (LMs) excel in language understanding and are mainly classified into two types: discriminative and generative. Discriminative models, like BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019) and SentenceBERT (Reimers and Gurevych, 2019), focus on estimating the conditional probability $P(\mathbf{y}|\mathbf{x})$ and are effective in tasks such as text classification and sentiment analysis. In contrast, generative models, including GPT-3 (Brown et al., 2020) and GPT-4 (OpenAI, 2024), aim to model the joint probability $P(\mathbf{x},\mathbf{y})$ for tasks like machine translation and text generation. These generative pre-trained models have significantly advanced the field of natural language processing (NLP) by leveraging massive datasets and billions of parameters, contributing to the rise of Large Language Models (LLMs) with outstanding performance across various tasks.
In the early stages, RAG and GraphRAG focused on improving pre-training techniques for discriminative language models (Devlin et al., 2019; Liu et al., 2019; Reimers and Gurevych, 2019). Recently, LLMs such as ChatGPT (Ouyang et al., 2022), LLaMA (Dubey et al., 2024), and Qwen2 (Yang et al., 2024c) have shown great potential in language understanding, demonstrating powerful in-context learning capabilities. Subsequently, research on RAG and GraphRAG shifted towards enhancing information retrieval for language models, addressing increasingly complex tasks and mitigating hallucinations, thereby driving rapid advancements in the field.
4. Overview of GraphRAG
GraphRAG is a framework that leverages external structured knowledge graphs to improve contextual understanding of LMs and generate more informed responses, as depicted in Figure 2. The goal of GraphRAG is to retrieve the most relevant knowledge from databases, thereby enhancing the answers of downstream tasks. The process can be defined as
(3) $a^{*}=\arg\max_{a\in A}p(a|q,\mathcal{G}),$
where $a^{*}$ is the optimal answer of the query $q$ given the TAG $\mathcal{G}$, and $A$ is the set of possible responses. After that, we jointly model the target distribution $p(a|q,\mathcal{G})$ with a graph retriever $p_{\theta}(G|q,\mathcal{G})$ and an answer generator $p_{\phi}(a|q,G)$ where $\theta,\phi$ are learnable parameters, and utilize the total probability formula to decompose $p(a|q,\mathcal{G})$, which can be formulated as
(4) $p(a|q,\mathcal{G})$ $=\sum\limits_{G\subseteq\mathcal{G}}p_{\phi}(a|q,G)p_{\theta}(G|q,\mathcal{G})$
$\approx p_{\phi}(a|q,G^{*})p_{\theta}(G^{*}|q,\mathcal{G}),$
where $G^{*}$ is the optimal subgraph. Because the number of candidate subgraphs can grow exponentially with the size of the graph, efficient approximation methods are necessary. The first line of Equation 4 is thus approximated by the second line. Specifically, a graph retriever is employed to extract the optimal subgraph $G^{*}$, after which the generator produces the answer based on the retrieved subgraph.
Therefore, in this survey, we decompose the entire process of GraphRAG into three main stages: Graph-Based Indexing, Graph-Guided Retrieval, and Graph-Enhanced Generation. The overall workflow of GraphRAG is illustrated in Figure 2 and detailed introductions of each stage are as follows.
Graph-Based Indexing (G-Indexing)
Graph-Based Indexing constitutes the initial phase of GraphRAG, aimed at identifying or constructing a graph database $\mathcal{G}$ that aligns with downstream tasks and establishing indices on it. The graph database can originate from public knowledge graphs (Vrandečić and Krötzsch, 2014; Bollacker et al., 2008a; Auer et al., 2007; Suchanek et al., 2007; Liu and Singh, 2004; Sap et al., 2019a), graph data (Morris et al., 2020), or be constructed based on proprietary data sources such as textual (Edge et al., 2024; Li et al., 2024e; Wang et al., 2024b; Gutiérrez et al., 2024) or other forms of data (Xu et al., 2024). The indexing process typically includes mapping node and edge properties, establishing pointers between connected nodes, and organizing data to support fast traversal and retrieval operations. Indexing determines the granularity of the subsequent retrieval stage, playing a crucial role in enhancing query efficiency.
Graph-Guided Retrieval (G-Retrieval)
Following graph-based indexing, the graph-guided retrieval phase focuses on extracting pertinent information from the graph database in response to user queries or input. Specifically, given a user query $q$ which is expressed in natural language, the retrieval stage aims to extract the most relevant elements (e.g., entities, triplets, paths, subgraphs) from knowledge graphs, which can be formulated as
[…]
$=\operatorname*{arg\,max}_{G\subseteq\mathcal{R}(\mathcal{G})}\,p_{\theta}(G|q,\mathcal{G})$
$=\operatorname*{arg\,max}_{G\subseteq\mathcal{R}(\mathcal{G})}\,\textbf{Sim}(q,G),$
where $G^{*}$ is the optimal retrieved graph elements and $\textbf{Sim}(\cdot,\cdot)$ is a function that measures the semantic similarity between user queries and the graph data. $\mathcal{R(\cdot)}$ represents a function to narrow down the search range of subgraphs, considering the efficiency.
Graph-Enhanced Generation (G-Generation)
The graph-enhanced generation phase involves synthesizing meaningful outputs or responses based on the retrieved graph data. This could encompass answering user queries, generating reports, etc. In this stage, a generator takes the query, retrieved graph elements, and an optional prompt as input to generate a response, which can be denoted as
[…]
$=\operatorname*{arg\,max}_{a\in A}p_{\phi}(a|q,G^{*})$
$=\operatorname*{arg\,max}_{a\in A}p_{\phi}(a|\mathcal{F}(q,G^{*})),$
where $\mathcal{F}(\cdot,\cdot)$ is a function that converts graph data into a form the generator can process.
5. Graph-Based Indexing
The construction and indexing of graph databases form the foundation of GraphRAG, where the quality of the graph database directly impacts GraphRAG’s performance. In this section, we categorize and summarize the selection or construction of graph data and various indexing methods that have been employed.
5.1. Graph Data
Various types of graph data are utilized in GraphRAG for retrieval and generation. Here, we categorize these data into two categories based on their sources, including Open Knowledge Graphs and Self-Constructed Graph Data.
5.1.1. Open Knowledge Graphs
Open knowledge graphs refer to graph data sourced from publicly available repositories or databases (Vrandečić and Krötzsch, 2014; Bollacker et al., 2008a; Auer et al., 2007; Suchanek et al., 2007). Using these knowledge graphs could dramatically reduce the time and resources required to develop and maintain. In this survey, we further classify them into two categories according to their scopes, i.e., General Knowledge Graphs and Domain Knowledge Graphs.
(1) General Knowledge Graphs
General knowledge graphs primarily store general, structured knowledge, and typically rely on collective input and updates from a global community, ensuring a comprehensive and continually refreshed repository of information.
Encyclopedic knowledge graphs are a typical type of general knowledge graph, which contains large-scale real-world knowledge collected from human experts and encyclopedias. For example, Wikidata [^1] https://www.wikidata.org/ (Vrandečić and Krötzsch, 2014) is a free and open knowledge base that stores structured data of its Wikimedia sister projects like Wikipedia, Wikivoyage, Wiktionary, and others. Freebase [^2] http://www.freebase.be/ (Bollacker et al., 2008a) is an extensive, collaboratively edited knowledge base that compiles data from various sources, including individual contributions and structured data from databases like Wikipedia. DBpedia [^3] https://www.dbpedia.org/ (Auer et al., 2007) represents information about millions of entities, including people, places, and things, by leveraging the infoboxes and categories present in Wikipedia articles. YAGO [^4] https://yago-knowledge.org/ (Suchanek et al., 2007) collects knowledge from Wikipedia, WordNet, and GeoNames.
Commonsense knowledge graphs are another type of general knowledge graph. They include abstract commonsense knowledge, such as semantic associations between concepts and causal relationships between events. Typical Commonsense Knowledge Graphs include: ConceptNet [^5] https://conceptnet.io/ (Liu and Singh, 2004) is a semantic network built from nodes representing words or phrases connected by edges denoting semantic relationships. ATOMIC (Sap et al., 2019a; Hwang et al., 2021) models the causal relationships between events.
(2) Domain Knowledge Graphs
As discussed in Section 1, domain-specific knowledge graphs are crucial for enhancing LLMs in addressing domain-specific questions. These KGs offer specialized knowledge in particular fields, aiding models in gaining deeper insights and a more comprehensive understanding of complex professional relationships. In the biomedical field, CMeKG [^6] https://cmekg.pcl.ac.cn/ encompasses a wide range of data, including diseases, symptoms, treatments, medications, and relationships between medical concepts. CPubMed-KG [^7] https://cpubmed.openi.org.cn/graph/wiki is a medical knowledge database in Chinese, building on the extensive repository of biomedical literature in PubMed. In the movie domain, Wiki-Movies (Miller et al., 2016) extracts structured information from Wikipedia articles related to films, compiling data about movies, actors, directors, genres, and other relevant details into a structured format. Additionally,  Jin et al. (2024b) construct a dataset named GR-Bench, which includes five domain knowledge graphs spanning academic, E-commerce, literature, healthcare, and legal fields. Furthermore,  He et al. (2024) convert triplet-format and JSON files from ExplaGraphs and SceneGraphs into a standard graph format and selects questions requiring 2-hop reasoning from WebQSP to create the universal graph-format dataset GraphQA for evaluating GraphRAG systems.
5.1.2. Self-Constructed Graph Data
Self-Constructed Graph Data facilitates the customization and integration of proprietary or domain-specific knowledge into the retrieval process. For downstream tasks that do not inherently involve graph data, researchers often propose constructing a graph from multiple sources (e.g., documents, tables, and other databases) and leveraging GraphRAG to enhance task performance. Generally, these self-constructed graphs are closely tied to the specific design of the method, distinguishing them from the open-domain graph data previously mentioned.
To model the structural relationships between the documents,  Munikoti et al. (2023) propose to construct a heterogeneous document graph capturing multiple document-level relations, including co-citation, co-topic, co-venue, etc.  Li et al. (2024b) and Wang et al. (2024b) establish relationship between passages according to shared keywords. To capture the relations between entities in documents,  Delile et al. (2024) , Edge et al. (2024) , Gutiérrez et al. (2024) and Li et al. (2024e) utilize the named entity recognition tools to extract entities from documents and language models to further extract relations between entities, where the retrieved entities and relations then form a knowledge graph. There are also some mapping methods for downstream tasks that need to be designed based on the characteristics of the task itself. For example, to solve the patent phrase similarity inference task,  Peng and Yang (2024) convert the patent database into a patent-phrase graph. Connections between patent nodes and phrase nodes are established if the phrases appear in the patents, while connections between patent nodes are based on citation relations. Targeting customer service technical support scenarios,  Xu et al. (2024) propose to model historical issues into a KG, which transforms the issues into tree representations to maintain the intra-issue relations, and utilize semantic similarities and a threshold to preserve inter-issue relations.
5.2. Indexing
Graph-Based Indexing plays a crucial role in enhancing the efficiency and speed of query operations on graph databases, directly influencing subsequent retrieval methods and granularity. Common graph-based indexing methods include graph indexing, text indexing, and vector indexing.
Figure 3. The overview of graph-based indexing.
5.2.1. Graph Indexing
Graph indexing represents the most commonly used approach, preserving the entire structure of the graph. This method ensures that for any given node, all its edges and neighboring nodes are easily accessible. During subsequent retrieval stages, classic graph search algorithms such as BFS and Shortest Path Algorithms can be employed to facilitate retrieval tasks (Jin et al., 2024b; Luo et al., 2024b; Sun et al., 2024b; Ma et al., 2024; Jiang et al., 2024a; Yasunaga et al., 2021; Taunk et al., 2023).
5.2.2. Text Indexing
Text indexing involves converting graph data into textual descriptions to optimize retrieval processes. These descriptions are stored in a text corpus, where various text-based retrieval techniques, such as sparse retrieval and dense retrieval, can be applied. Some approaches transform knowledge graphs into human-readable text using predefined rules or templates. For instance, Li et al. (2023) , Huang et al. (2023a) and Li et al. (2024a) use predefined templates to convert each triple in knowledge graphs into natural language, while Yu et al. (2023) merge triplets with the same head entity into passages. Additionally, some methods convert subgraph-level information into textual descriptions. For example, Edge et al. (2024) perform community detection on the graph and generate summaries for each community using LLMs.
5.2.3. Vector Indexing
Vector indexing transforms graph data into vector representations to enhance retrieval efficiency, facilitating rapid retrieval and effective query processing. For example, entity linking can be seamlessly applied through query embeddings, and efficient vector search algorithms such as Locality Sensitive Hashing (LSH) (Jafari et al., 2021) can be utilized. G-Retriever (He et al., 2024) employs language models to encode textual information associated with each node and edge within the graph, while GRAG (Hu et al., 2024) uses language models to convert $k$-hop ego networks into graph embeddings, thereby better preserving structural information.
5.2.4. Hybrid Indexing
Each of the above three indexing methods offers distinct advantages: graph indexing facilitates easy access to structural information, text indexing simplifies retrieval of textual content, and vector indexing enables quick and efficient searches. Therefore, in practical applications, a hybrid approach combining these indexing methods is often preferred over relying solely on one. For instance, HybridRAG (Sarmah et al., 2024) retrieves both vector and graph data simultaneously to enhance the content retrieved. While EWEK-QA (Dehghan et al., 2024) uses both text and knowledge graphs.
Figure 4. The general architectures of graph-based retrieval.
6. Graph-Guided Retrieval
In GraphRAG, the retrieval process is crucial for ensuring the quality and relevance of generated outputs by extracting pertinent and high-quality graph data from external graph databases. However, retrieving graph data presents two significant challenges: (1) Explosive Candidate Subgraphs: As the graph size increases, the number of candidate subgraphs grows exponentially, requiring heuristic search algorithms to efficiently explore and retrieve relevant subgraphs. (2) Insufficient Similarity Measurement: Accurately measuring similarity between textual queries and graph data necessitates the development of algorithms capable of understanding both textual and structural information.
Considerable efforts have previously been dedicated to optimizing the retrieval process to address the above challenges. This survey focuses on examining various aspects of the retrieval process within GraphRAG, including the selection of the retriever, retrieval paradigm, retrieval granularity, and effective enhancement techniques. The general architectures of Graph-Guided Retrieval are depicted in Figure 4.
6.1. Retriever
In GraphRAG, various retrievers possess unique strengths for addressing different aspects of retrieval tasks. We categorize retrievers into three types based on their underlying models: Non-parametric Retriever, LM-based Retriever, and GNN-based Retriever. It is important to note that models used in pre-processing steps, such as query encoding and entity linking, are not considered here, as these models vary across different methods and are not the primary focus of this paper.
6.1.1. Non-parametric Retriever
Non-parametric retrievers, based on heuristic rules or traditional graph search algorithms, do not rely on deep-learning models, thereby achieving high retrieval efficiency. For instance, Yasunaga et al. (2021) and Taunk et al. (2023) retrieve $k$-hop paths containing the topic entities of each question-choice pair. G-Retriever (He et al., 2024) enhances the conventional Prize-Collecting Steiner Tree (PCST) algorithm by incorporating edge prices and optimizing relevant subgraph extraction. Delile et al. (2024) and Mavromatis and Karypis (2024) first extract entities mentioned in the query and then retrieve the shortest path related to these entities. These methods often involve an entity linking pre-processing step to identify nodes in the graph before retrieval.
6.1.2. LM-based Retriever
LMs serve as effective retrievers in GraphRAG due to their strong natural language understanding capabilities. These models excel in processing and interpreting diverse natural language queries, making them versatile for a wide range of retrieval tasks within graph-based frameworks. We primarily categorized LMs into two types: discriminative and generative language models. Subgraph Retriever (Zhang et al., 2022b) trains RoBERTa (Liu et al., 2019) as the retriever, which expands from the topic entity and retrieves the relevant paths in a sequential decision process. KG-GPT (Kim et al., 2023a) adopts LLMs to generate the set of top-$K$ relevant relations of the specific entity.  Wold et al. (2023) utilize fine-tuned GPT-2 to generate reasoning paths. StructGPT (Jiang et al., 2023a) utilizes LLMs to automatically invoke several pre-defined functions, by which relevant information can be retrieved and combined to assist further reasoning.
6.1.3. GNN-based Retriever
GNNs are adept at understanding and leveraging complex graph structures. GNN-based retrievers typically encode graph data and subsequently score different retrieval granularities based on their similarity to the query. For example, GNN-RAG (Mavromatis and Karypis, 2024) first encodes the graph, assigns a score to each entity, and retrieves entities relevant to the query based on a threshold. EtD (Liu et al., 2024e) iterates multiple times to retrieve relevant paths. During each iteration, it first uses LLaMA2 (Touvron et al., 2023) to select edges connecting the current node, then employs GNNs to obtain embeddings of the new layer of nodes for the next round of LLM selection.
6.1.4. Discussion
During the retrieval process, non-parametric retrievers exhibit good retrieval efficiency, but they may suffer from inaccurate retrieval due to a lack of training on downstream tasks. Meanwhile, although LM-based retrievers and GNN-based retrievers offer higher retrieval accuracy, they require significant computational overhead. Considering this complementarity, many methods propose hybrid retrieval approaches to improve both retrieval efficiency and accuracy. Many approaches adopt a multi-stage retrieval strategy, employing different models at each stage. For example, RoG (Luo et al., 2024b) first utilizes LLMs to generate planning paths and then extracts paths satisfying the planning paths from knowledge graphs. GenTKGQA (Gao et al., 2024a) infers crucial relations and constraints from the query using LLMs and extracts triplets according to these constraints.
6.2. Retrieval Paradigm
Within GraphRAG, different retrieval paradigms, including once retrieval, iterative retrieval, and multi-stage retrieval, play crucial roles in improving the relevance and depth of the retrieved information. Once retrieval aims to gather all pertinent information in a single operation. Iterative retrieval conducts further searches based on previously retrieved information, progressively narrowing down to the most relevant results. Here we further divide iterative retrieval into adaptive retrieval and non-adaptive retrieval, with the only difference lying in whether the stopping of the retrieval is determined by the model. Another retrieval paradigm is multi-stage retrieval, where retrieval is divided into multiple stages. Different types of retrievers may be employed at each stage for more precise and diversified search results. Below, we will provide a detailed introduction to these types of retrieval paradigms.
6.2.1. Once Retrieval
Once retrieval aims to retrieve all the relevant information in a single query. One category of approaches (Gutiérrez et al., 2024; Li et al., 2023; Hu et al., 2024) utilize embedding similarities to retrieve the most relevant pieces of information. Another category of methods design pre-defined rules or patterns to directly extract specific structured information such as triplets, paths or subgraphs from graph databases. For example, G-Retriever (He et al., 2024) utilizes an extended PCST algorithm to retrieve the most relevant subgraph. KagNet (Lin et al., 2019) extracts paths between all pairs of topic entities with lengths not exceeding $k$Yasunaga et al. (2021) and Taunk et al. (2023) extract the subgraph that contains all topic entities along with their $2$-hop neighbors.
Furthermore, in this subsection, we also include some multiple retrieval methods that involve decoupled and independent retrievals, allowing them to be computed in parallel and executed only once. For example, Luo et al. (2024b) and Cheng et al. (2024) first instruct LLMs to generate multiple reasoning paths and then use a BFS retriever to sequentially search for subgraphs in the knowledge graphs that match each path. KG-GPT (Kim et al., 2023a) decomposes the original query into several sub-queries, retrieving relevant information for each sub-query in a single retrieval process.
6.2.2. Iterative Retrieval
In iterative retrieval, multiple retrieval steps are employed, with subsequent searches depending on the results of prior retrievals. These methods aim to deepen the understanding or completeness of the retrieved information over successive iterations. In this survey, we further classify iterative retrieval into two categories: (1) non-adaptive and (2) adaptive retrieval. We provide a detailed summary of these two categories of methods below.
(1) Non-Adaptive Retrieval
Non-adaptive methods typically follow a fixed sequence of retrieval, and the termination of retrieval is determined by setting a maximum time or a threshold. For example, PullNet (Sun et al., 2019) retrieves problem-relevant subgraphs through $T$ iterations. In each iteration, the paper designs a retrieval rule to select a subset of retrieved entities, and then expands these entities by searching relevant edges in the knowledge graph. In each iteration, KGP (Wang et al., 2024b) first selects seed nodes based on the similarity between the context and the nodes in the graph. It then uses LLMs to summarize and update the context of the neighboring nodes of the seed nodes, which is utilized in the subsequent iteration.
(2) Adaptive Retrieval
One distinctive characteristic of adaptive retrieval is to let models autonomously determine the optimal moments to finish the retrieval activities. For instance, (Wu et al., 2023b; Guo et al., 2024) leverage an LM for hop prediction, which serves as an indicator to end the retrieval. There is also a group of researchers who utilize model-generated special tokens or texts as termination signals for the retrieval process. For example, ToG (Sun et al., 2024b; Ma et al., 2024) prompts the LLM agent to explore the multiple possible reasoning paths until the LLM determines the question can be answered based on the current reasoning path. (Zhang et al., 2022b) trains a RoBERTa to expand a path from each topic entity. In the process, a virtual relation named as “[END]” is introduced to terminate the retrieval process.
Another common approach involves treating the large model as an agent, enabling it to directly generate answers to questions to signal the end of iteration. For instance, (Jin et al., 2024b; Jiang et al., 2023a, 2024b; Wang et al., 2023d; Sun et al., 2024a) propose LLM-based agents to reason on graphs. These agents could autonomously determine the information for retrieval, invoke the pre-defined retrieval tools, and cease the retrieval process based on the retrieved information.
6.2.3. Multi-Stage Retrieval
Multi-stage retrieval divides the retrieval process linearly into multiple stages, with additional steps such as retrieval enhancement, and even generation processes occurring between these stages. In multi-stage retrieval, different stages may employ various types of retrievers, which enables the system to incorporate various retrieval techniques tailored to different aspects of the query. For example,  Wang et al. (2024a) first utilize a non-parametric retriever to extract $n$-hop paths of entities in the query’s reasoning chain, then after a pruning stage, it further retrieves the one-hop neighbors of the entities in the pruned subgraph. OpenCSR (Han et al., 2023) divides the retrieval process into two stages. In the first stage, it retrieves all $1$-hop neighbors of the topic entity. In the second stage, it compares the similarity between these neighbor nodes and other nodes, selecting the top-$k$ nodes with the highest similarity for retrieval. GNN-RAG (Mavromatis and Karypis, 2024) first employs GNNs to retrieve the top-$k$ nodes most likely to be the answer. Subsequently, it retrieves all shortest paths between query entities and answer entities pairwise.
6.2.4. Discussion
In GraphRAG, once retrieval typically exhibits lower complexity and shorter response times, making it suitable for scenarios requiring real-time responsiveness. In contrast, iterative retrieval often involves higher time complexity, especially when employing LLMs as retrievers, potentially leading to longer processing times. However, this approach can yield higher retrieval accuracy by iteratively refining retrieved information and generating responses. Therefore, the choice of retrieval paradigm should balance accuracy and time complexity based on specific use cases and requirements.
6.3. Retrieval Granularity
According to different task scenarios and indexing types, researchers design distinct retrieval granularities (i.e., the form of related knowledge retrieved from graph data), which can be divided into nodes, triplets, paths, and subgraphs. Each retrieval granularity has its own advantages, making it suitable for different practical scenarios. We will introduce the details of these granularities in the following sections.
6.3.1. Nodes
Nodes allow for precise retrieval focused on individual elements within the graph, which is ideal for targeted queries and specific information extraction. In general, for knowledge graphs, nodes refer to entities. For other types of text attribute graphs, nodes may include textual information that describes the node’s attributes. By retrieving nodes within the graph, GraphRAG systems could provide detailed insights into their attributes, relationships, and contextual information. For example,  Munikoti et al. (2023) , Li et al. (2024b) and Wang et al. (2024b) construct document graphs and retrieves relevant passage nodes.  Liu et al. (2024e) , Sun et al. (2019) and Gutiérrez et al. (2024) retrieve entities from constructed knowledge graphs.
6.3.2. Triplets
Generally, triplets consist of entities and their relationships in the form of subject-predicate-object tuples, providing a structured representation of relational data within a graph. The structured format of triplets allows for clear and organized data retrieval, making it advantageous in scenarios where understanding relationships and contextual relevance between entities is critical.  Yang et al. (2024a) retrieve triplets containing topic entities as relevant information.  Huang et al. (2023a) , Li et al. (2023) and Li et al. (2024a) first convert each triplet of graph data into textual sentences using predefined templates and subsequently adopt a text retriever to extract relevant triplets. However, directly retrieving triplets from graph data may still lack contextual breadth and depth, thus being unable to capture indirect relationships or reasoning chains. To address this challenge,  Wang et al. (2023c) propose to generate the logical chains based on the original question, and retrieve the relevant triplets of each logical chain.
6.3.3. Paths
The retrieval of path-granularity data can be seen as capturing sequences of relationships between entities, enhancing contextual understanding and reasoning capabilities. In GraphRAG, retrieving paths offers distinct advantages due to their ability to capture complex relationships and contextual dependencies within a graph.
However, path retrieval can be challenging due to the exponential growth in possible paths as graph size increases, which escalates computational complexity. To address this, some methods retrieve relevant paths based on pre-defined rules. For example,  Wang et al. (2024a) and Lo and Lim (2023) first select entity pairs in the query and then traverse to find all the paths between them within $n$-hop. HyKGE (Jiang et al., 2024a) first defines three types of paths: path, co-ancestor chain, and co-occurrence chain, and then utilizes corresponding rules to retrieve each of these three types of paths. In addition, some methods utilize models to perform path searching on graphs. ToG (Sun et al., 2024b; Ma et al., 2024) proposes to prompt the LLM agent to perform the beam search on KGs and find multiple possible reasoning paths that help answer the question.  Luo et al. (2024b) , Wu et al. (2023b) and Guo et al. (2024) first utilizes the model to generate faithful reasoning plans and then retrieves relevant paths based on these plans. GNN-RAG (Mavromatis and Karypis, 2024) first identifies the entities in the question. Subsequently, all paths between entities that satisfy a certain length relationship are extracted.
6.3.4. Subgraphs
Retrieving subgraphs offers significant advantages due to its ability to capture comprehensive relational contexts within a graph. This granularity enables GraphRAG to extract and analyze complex patterns, sequences, and dependencies embedded within larger structures, facilitating deeper insights and a more nuanced understanding of semantic connections.
To ensure both information completeness and retrieval efficiency, some methods propose an initial rule-based approach to retrieve candidate subgraphs, which are subsequently refined or processed further.  Peng and Yang (2024) retrieve the ego graph of the patent phrase from the self-constructed patent-phrase graph.  Yasunaga et al. (2021) , Feng et al. (2020) and Taunk et al. (2023) first select the topic entities and their two-hop neighbors as the node set, and then choose the edges with head and tail entities both in the node set to form the subgraph. Besides, there are also some embedding-based subgraph retrieval methods. For example,  Hu et al. (2024) first encode all the $k$-hop ego networks from the graph database, then retrieve subgraphs related to the query based on the similarities between embeddings.  Wen et al. (2024) and Li et al. (2024e) extract two types of graphs, including Path evidence subgraphs and Neighbor evidence subgraphs, based on pre-defined rules. OpenCSR (Han et al., 2023) starts from a few initial seed nodes and gradually expands to new nodes, eventually forming a subgraph.
In addition to the aforementioned direct subgraph retrieval methods, some works propose first retrieving relevant paths and then constructing related subgraphs from them. For instance, Zhang et al. (2022b) train a RoBERTa model to identify multiple reasoning paths through a sequential decision process, subsequently merging identical entities from different paths to induce a final subgraph.
6.3.5. Hybrid Granularties
Considering the advantages and disadvantages of various retrieval granularities mentioned above, some researchers propose using hybrid granularities, that is, retrieving relevant information of multiple granularities from graph data. This type of granularity enhances the system’s ability to capture both detailed relationships and broader contextual understanding, thus reducing noise while improving the relevance of the retrieved data. Various previous works propose to utilize LLM agents to retrieve complex hybrid information.  Jin et al. (2024b) , Jiang et al. (2023a) , Jiang et al. (2024b) , Wang et al. (2023d) and Sun et al. (2024a) propose to adopt LLM-based agents for adaptively selecting nodes, triplets, paths, and subgraphs.
6.3.6. Discussion
(1) In real applications, there are no clear boundaries between these retrieval granularities, as subgraphs can be composed of multiple paths, and paths can be formed by several triplets. (2) Various granularities such as nodes, triplets, paths, and subgraphs offer distinct advantages in the GraphRAG process. Balancing between retrieval content and efficiency is crucial when selecting the granularity, depending on the specific context of the task. For straightforward queries or when efficiency is paramount, finer granularities such as entities or triplets may be preferred to optimize retrieval speed and relevance. In contrast, complex scenarios often benefit from a hybrid approach that combines multiple granularities. This approach ensures a more comprehensive understanding of the graph structure and relationships, enhancing the depth and accuracy of the generated responses. Thus, GraphRAG’s flexibility in granularity selection allows it to adapt effectively to diverse information retrieval needs across various domains.
6.4. Retrieval Enhancement
To ensure high retrieval quality, researchers propose techniques to enhance both user queries and the knowledge retrieved. In this paper, we categorize query enhancement into query expansion and query decomposition, and knowledge enhancement into merging and pruning. These strategies collectively optimize the retrieval process. Although other techniques such as query rewriting (Ma et al., 2023; Peng et al., 2024; Mao et al., 2024b; Qiao et al., 2024) are commonly used in RAG, they are less frequently applied in GraphRAG. We do not delve into these methods, despite their potential adaptation for GraphRAG.
6.4.1. Query Enhancement
Strategies applied to queries typically involve pre-processing techniques that enrich the information for better retrieval. This may include query expansion and query decomposition.
(1) Query Expansion
Due to the generally short length of queries and their limited information content, query expansion aims to improve search results by supplementing or refining the original query with additional relevant terms or concepts.  Luo et al. (2024b) generate relation paths grounded by KGs with LLMs to enhance the retrieval query.  Cheng et al. (2024) adopt SPARQL to get all the aliases of the query entities from Wikidata to augment the retrieval queries, which capture lexical variations of the same entity.  Huang et al. (2023a) propose a consensus-view knowledge retrieval method to improve retrieval accuracy, which first discover semantically relevant queries, and then re-weight the original query terms to enhance the retrieval performance. HyKGE (Jiang et al., 2024a) utilizes a large model to generate the hypothesis output of the question, concatenating the hypothesis output with the query as input to the retriever. Golden-Retriever (An et al., 2024) first recognizes the jargon in the query and then retrieves explanations of the jargon as a supplement to the query.
(2) Query Decomposition
Query decomposition techniques break down or decompose the original user query into smaller, more specific sub-queries. Each sub-query typically focuses on a particular aspect or component of the original query, which successfully alleviates the complexity and ambiguity of language queries. For instance, (Kim et al., 2023a; Choudhary and Reddy, 2024) breaks down the primary question into sub-sentences, each representing a distinct relation, and sequentially retrieves the pertinent triplets for each sub-sentence.
6.4.2. Knowledge Enhancement
After retrieving initial results, knowledge enhancement strategies are employed to refine and improve the retriever’s results. This phase often involves knowledge merging and knowledge pruning processes to present the most pertinent information prominently. These techniques aim to ensure that the final set of retrieved results is not only comprehensive but also highly relevant to the user’s information needs.
(1) Knowledge Merging
Knowledge merging retrieved information enables compression and aggregation of information, which assists in obtaining a more comprehensive view by consolidating relevant details from multiple sources. This approach not only enhances the completeness and coherence of the information but also mitigates issues related to input length constraints in models. KnowledgeNavigator (Guo et al., 2024) merges nodes and condenses the retrieved sub-graph through triple aggregation to enhance the reasoning efficiency. In Subgraph Retrieval (Zhang et al., 2022b), after retrieving top-$k$ paths from each topic entity to form a single subgraph, researchers propose to merge the same entities from different subgraphs to form the final subgraph.  Wen et al. (2024) and Li et al. (2024e) merge retrieved subgraphs based on relations, combining head entities and tail entities that satisfy the same relation into two distinct entity sets, ultimately forming a relation paths.
(2) Knowledge Pruning
Knowledge pruning involves filtering out less relevant or redundant retrieved information to refine the results. Previous approaches for pruning encompass two main categories: (re)-ranking-based approaches and LLM-based approaches. (Re)-ranking methods involve the reordering or prioritization of retrieved information using tailored metrics or criteria.
One line of methods introduces stronger models for reranking. For example,  Li et al. (2023) concatenate each retrieved triplet with the question-choice pair, and adopt a pre-trained cross-encoder (Reimers and Gurevych, 2019) to re-rank the retrieved triplets.  Jiang et al. (2024a) utilize the FlagEmbedding to encode the text to re-rank top-k documents returned by embedding model “bge_reranker_large”.  Liu et al. (2024b) train a PLM to
Another category utilizes the similarity between queries and retrieved information for ranking. For instance,  Cheng et al. (2024) re-rank the candidate subgraphs based on the similarity for both relation and fine-grained concept between subgraphs and the query.  Taunk et al. (2023) first cluster the 2-hop neighbors and then delete the cluster with the lowest similarity score with the input query.  Yasunaga et al. (2021) prune the retrieved subgraph according to the relevance score between the question context and the KG entity nodes calculated by a pre-trained language model.  Wang et al. (2024a) , Jiang et al. (2023b) , Gutiérrez et al. (2024) and Luo et al. (2023) adopt Personalized PageRank algorithm to rank the retrieved candidate information for further filtering.  Liu et al. (2024b) trains a PLM to score the similarity between the retrieved information and the query, and rerank the retrieved paths based on the similarity score. G-G-E (Gao et al., 2022) first divides the retrieved subgraph into several smaller subgraphs, then compares the similarity between each smaller subgraph and the query. Subgraphs with low similarity are removed, and the remaining smaller subgraphs are merged into a larger subgraph.
Additionally, a third category of methods proposes new metrics for reranking. For example,   Munikoti et al. (2023) propose a metric that measures both the impact and recency of the retrieved text chunks. KagNet (Lin et al., 2019) decomposes the retrieved paths into triplets and reranks the paths based on the confidence score measured by the knowledge graph embedding (KGE) techniques. LLM-based methods excel in capturing complex linguistic patterns and semantic nuances, which enhances their ability to rank search results or generate responses more accurately. To avoid introducing noisy information,   Wang et al. (2024a) and Kim et al. (2023a) propose to prune the irrelevant graph data by calling LLMs to check.
Figure 5. The overview of graph-enhanced generation.
7. Graph-Enhanced Generation
The generation stage is another crucial step in GraphRAG, aimed at integrating the retrieved graph data with the query to enhance response quality. In this stage, suitable generation models must be selected based on the downstream tasks. The retrieved graph data is then transformed into formats compatible with the generators. The generator takes both the query and the transformed graph data as inputs to produce the final response. Beyond these fundamental processes, generative enhancement techniques can further improve the output by intensifying the interaction between the query and the graph data and enriching the content generation itself. The organization of this section and the overview of graph-enhanced generation are depicted in Figure 5.
7.1. Generators
The selection of generators often depends on the type of downstream task at hand. For discriminative tasks (e.g., multi-choice question answering) or generative tasks that can be formulated as discriminative tasks (e.g., KBQA), one can utilize GNNs or discriminative language models to learn representations of the data. These representations can then be mapped to the logits associated with different answer options to provide responses. Alternatively, generative language models can be employed to directly generate answers. For generative tasks, however, the use of GNNs and discriminative language models alone is insufficient. These tasks require the generation of text, which necessitates the deployment of decoders.
7.1.1. GNNs
Due to the powerful representational capabilities of GNNs for graph data, they are particularly effective for discriminative tasks. GNNs can directly encode graph data, capturing complex relationships and node features inherent in the graph structure. This encoding is then processed through a Multi-Layer Perceptron (MLP) to generate predictive outcomes. These approaches primarily utilize classical GNN models (e.g., GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), GraphSAGE (Hamilton et al., 2017), and Graph Transformers (Shehzad et al., 2024)), either in their original form or modified to better align with downstream tasks. For example, HamQA (Dong et al., 2023b) designs a hyperbolic GNN to learn the representations of retrieved graph data, which learns from the mutual hierarchical information between query and graphs.  Sun et al. (2018) compute PageRank scores for neighboring nodes and aggregates them weighted by these scores, during message-passing. This approach enhances the central node’s ability to assimilate information from its most relevant neighboring nodes.  Mavromatis and Karypis (2022) decode the query into several vectors (instructions), and enhances instruction decoding and execution for effective reasoning by emulating breadth-first search (BFS) with GNNs to improve instruction execution and using adaptive reasoning to update the instructions with KG-aware information.
7.1.2. LMs
LMs possess strong capabilities in text understanding, which also allows them to function as generators. In the context of integrating LMs with graph data, it is necessary to first convert the retrieved graph data into specific graph formats. This conversion process ensures that the structured information is effectively understood and utilized by the LMs. These formats, which will be elaborated on in Section 7.2, are crucial for preserving the relational and hierarchical structure of the graph data, thereby enhancing the model’s ability to interpret complex data types. Once the graph data is formatted, it is then combined with a query and fed into an LM.
For encoder-only models, such as BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019), their primary use is in discriminative tasks. Similar to GNNs, these models first encode the input text and then utilize MLPs to map it to the answer space (Jiang et al., 2023b; Huang et al., 2023a; Li et al., 2023). On the other hand, encoder-decoder and decoder-only models, such as T5 (Raffel et al., 2020), GPT-4 (OpenAI, 2024), and LLaMA (Dubey et al., 2024), are adept at both discriminative and generative tasks. These models excel in text understanding, generation, and reasoning, allowing them to process textual inputs directly and generate textual responses (Edge et al., 2024; Wang et al., 2024a; Luo et al., 2024b; Jin et al., 2024b; Sun et al., 2024b; Mavromatis and Karypis, 2024; Wang et al., 2023c; Jiang et al., 2024a).
7.1.3. Hybrid Models
Considering the strengths of GNNs at representing the structure of graph data, and the robust understanding of text demonstrated by LMs, many studies are exploring the integration of these two technologies to generate coherent responses. This paper categorizes the hybrid generative approaches into two distinct types: cascaded paradigm and parallel paradigm.
(1) Cascaded Paradigm
In the cascaded approaches, the process involves a sequential interaction where the output from one model serves as the input for the next. Specifically, the GNN processes the graph data first, encapsulating its structural and relational information into a form that the LM can understand. Subsequently, this transformed data is fed into the LM, which then generates the final text-based response. These methods leverage the strengths of each model in a step-wise fashion, ensuring detailed attention to both structural and textual data.
In these methods, prompt tuning (Liu et al., 2023, 2022; Li and Liang, 2021; Lester et al., 2021) is a typical approach, where GNNs are commonly employed to encode the retrieved graph data. The encoded graph data is subsequently pre-pended as a prefix to the input text embeddings of an LM. The GNN is then optimized through downstream tasks to produce enhanced encodings of the graph data (He et al., 2024; Hu et al., 2024; Zhang et al., 2024b; Gao et al., 2024a).
(2) Parallel Paradigm
On the other hand, the parallel approach operates by concurrently utilizing the capabilities of both the GNN and the LLM. In this setup, both models receive the initial inputs simultaneously and work in tandem to process different facets of the same data. The outputs are then merged, often through another model or a set of rules, to produce a unified response that integrates insights from both the graphical structure and the textual content.
In the parallel paradigm, a typical approach involves separately encoding inputs using both GNNs and LMs, followed by integrating these two representations, or directly integrating their output responses. For instance,  Jiang et al. (2022) aggregate predictions from GNNs and LMs by weighted summation to obtain the final answer.  Lin et al. (2019) and Pahuja et al. (2023) integrate the graph representations derived from GNNs and the text representations generated by LMs using attention mechanisms. Yasunaga et al. (2021) , Munikoti et al. (2023) and Taunk et al. (2023) directly concatenate graph representations with text representations.
Another approach involves designing dedicated modules that integrate GNNs with LMs, enabling the resulting representations to encapsulate both structural and textual information. For instance, Zhang et al. (2022a) introduce a module called the GreaseLM Layer, which incorporates both GNN and LM layers. At each layer, this module integrates textual and graph representations using a two-layer MLP before passing them to the next layer. Similarly, ENGINE (Zhu et al., 2024) proposes G-Ladders, which combine LMs and GNNs through a side structure, enhancing node representations for downstream tasks.
Discussion
Hybrid models that harness both the representation capabilities of GNNs for graph data and LMs for text data hold promising applications. However, effectively integrating information from these two modalities remains a significant challenge.
7.2. Graph Formats
When using GNNs as generators, the graph data can be directly encoded. However, when utilizing LMs as generators, the non-Euclidean nature of graph data poses a challenge, as it cannot be directly combined with textual data for input into the LMs. To address this, graph translators are employed to convert the graph data into a format compatible with LMs. This conversion enhances the generative capabilities of LMs by enabling them to effectively process and utilize structured graph information. In this survey, we summarize two distinct graph formats: graph languages and graph embeddings. We illustrate this process with an example in Figure 6, with detailed introductions provided below.
Figure 6. Illustration of the graph languages. Given the retrieved subgraph on the left part, we show how to transform it into adjacency/edge table, natural language, node sequence, code-like forms and syntax trees to adapt the input form requirements of different generators.
7.2.1. Graph Languages
A graph description language is a formalized system of notation that is specifically crafted to characterize and represent graph data. It prescribes a uniform syntax and semantic framework that describes the components and interconnections within a graph. Through these languages, users can consistently generate, manipulate, and interpret graph data in a comprehensible format to machines. They enable the definition of graph architectures, the specification of attributes for nodes and edges, and the implementation of operations and queries on graph structures. Next, we will introduce five types of graph languages separately: Adjacency / Edge Table, Natural Language, Codes, Syntax Tree, and Node Sequence.
(1) Adjacency / Edge Table
The adjacency table and the edge table are widely used methods for describing graph structures (Fatemi et al., 2023; Guo et al., 2023; Wang et al., 2023b; Li et al., 2024f). The adjacency table enumerates the immediate neighbors of each vertex, offering a compact way to represent connections in sparse graphs. For example, KG-GPT (Kim et al., 2023a) linearizes the triples in the retrieved subgraph, which are then concatenated and fed into the LLMs. Conversely, the edge table details all the edges within the graph, providing a straightforward representation that is particularly useful for processing and analyzing graphs in a linear format. Both two methods are brief, easy to understand, and intuitive.
(2) Natural Language
Given that user queries are typically presented in natural language, and considering the outstanding natural language comprehension capabilities of LMs, it becomes a compelling approach to describe the retrieved graph data using natural language. By translating graph data into descriptive, easily comprehensible language, LMs can bridge the gap between raw data representation and user-friendly information, facilitating more effective interactions with data-driven applications. For example, some researchers (Li et al., 2023; Huang et al., 2023a) propose defining a natural language template for each type of edge in advance and subsequently filling in the endpoints of each edge into the corresponding template based on its type.  Ye et al. (2024) employ natural language to describe the information of $1$-hop and $2$-hop neighboring nodes of the central node.  Edge et al. (2024) utilize LLMs to generate report-like summaries for each detected graph community.  Wu et al. (2023b) and Guo et al. (2024) adopt LMs to rewrite the edge table of retrieved subgraphs, generating a natural language description.  Fatemi et al. (2023) explore different representations of nodes (e.g., Integer encoding, alphabet letters, names, etc.) and edges (e.g., parenthesis, arrows, incident, etc.).  Jin et al. (2024b) , Jiang et al. (2023a) , Jiang et al. (2024b) , Wang et al. (2023d) and Sun et al. (2024a) integrate information from different granularities within the graph into prompts through natural language in the form of dialogue.
(3) Code-Like Forms
Considering that natural language descriptions and other $1$-D sequences are inherently inadequate for directly representing the $2$-D structure of graph data, and given the robust code comprehension capabilities of LMs, many researchers (Guo et al., 2023) explore using code-like formats to represent graph structures. For example,   Guo et al. (2023) examine the use of Graph Modeling Language (GML) (Himsolt, 1996) and Graph Markup Language (GraphML) (Rong et al., 2020) for representing graphs. These standardized languages are specifically designed for graph data, providing comprehensive descriptions that encompass nodes, edges, and their interrelationships.
(4) Syntax Tree
Compared to direct flattening of graphs, some research (Zhao et al., 2023) propose transforming graphs into structures akin to syntax trees. Syntax trees possess a hierarchical structure and, being topological graphs, also maintain a topological order. This method retains more structural information, enhancing the understanding and analysis of the graph’s intrinsic properties. Such a transformation not only preserves the relational dynamics between different graph elements but also facilitates more sophisticated algorithms for graph analysis and processing. GRAPHTEXT (Zhao et al., 2023) proposes transforming the ego network of a central node into a graph-syntax tree format. This format not only encapsulates structural information but also integrates the features of the nodes. By traversing this syntax tree, it is possible to obtain a node sequence that maintains both topological order and hierarchical structure.
(5) Node Sequence
Some studies (Chen et al., 2024; Mavromatis and Karypis, 2024) propose representing graphs through sequences of nodes, which are often generated using predefined rules. Compared to natural language descriptions, these node sequences are more concise and incorporate prior knowledge, specifically the structural information emphasized by the rules.  Luo et al. (2024b) and Sun et al. (2024b) transform the retrieved paths into node sequences and input them into an LLM to enhance the task performance. LLaGA (Chen et al., 2024) proposes two templates that can transform graphs into node sequences. The first template, known as the Neighborhood Detail Template, offers a detailed examination of the central node along with its immediate surroundings. The second, termed the Hop-Field Overview Template, provides a summarized perspective of a node’s neighborhood, which can be expanded to encompass broader areas. GNN-RAG (Mavromatis and Karypis, 2024) inputs the retrieved reasoning paths into LMs in the form of node sequences as prompts.
Discussion
Good graph languages should be complete, concise, and comprehensible. Completeness entails capturing all essential information within the graph structure, ensuring no critical details are omitted. Conciseness refers to the necessity of keeping textual descriptions brief to avoid the “lost in the middle” phenomenon (Liu et al., 2024a) or exceeding the length limitations of LMs. Lengthy inputs can hinder LMs’ processing capabilities, potentially causing loss of context or truncated data interpretation. Comprehensibility ensures that the language used is easily understood by LLMs, facilitating accurate representation of the graph’s structure. Due to the characteristics of different graph languages, their choice can significantly impact the performance of downstream tasks (Fatemi et al., 2023).
7.2.2. Graph Embeddings
The above graph language methods transform graph data into text sequences, which may result in overly lengthy contexts, incurring high computational costs and potentially exceeding the processing limits of LLMs. Additionally, LLMs currently struggle to fully comprehend graph structures even with graph languages (Guo et al., 2023). Thus, using GNNs to represent graphs as embeddings presents a promising alternative. The core challenge lies in integrating graph embeddings with textual representations into a unified semantic space. Current research focuses on utilizing prompt tuning methodologies, as discussed earlier. There are also some methods that adopt FiD (Fusion-in-Decoder) (Izacard and Grave, 2021; Yu et al., 2022), which first convert the graph data into text, then encode it using an LM-based encoder and input it into the decoders (Fang et al., 2024a; Yu et al., 2023; Dong et al., 2023a). Notably, feeding graph representations into LMs is feasible primarily with open-source LMs, not closed-source models like GPT-4 (OpenAI, 2024). While graph embedding methods avoid handling long text inputs, they face other challenges, such as difficulty in preserving precise information like specific entity names and poor generalization.
7.3. Generation Enhancement
In the generation phase, besides converting the retrieved graph data into formats acceptable by the generator and inputting it together with the query to generate the final response, many researchers explore various methods of generation enhancement techniques to improve the quality of output responses. These methods can be classified into three categories based on their application stages: pre-generation enhancement, mid-generation enhancement, and post-generation enhancement.
7.3.1. Pre-Generation Enhancement
Pre-generation enhancement techniques focus on improving the quality of input data or representations before feeding them into the generator. In fact, there is no clear boundary between Pre-Generation Enhancement and Retrieval. In this survey, we categorize the retrieval stage as the process of retrieving knowledge from the original graph, and merging and pruning retrieved knowledge. Subsequent operations are considered Pre-Generation Enhancements.
Commonly used pre-generation enhancement approaches primarily involve semantically enriching the retrieved graph data to achieve tighter integration between the graph data and textual query.  Wu et al. (2023b) employ LLMs to rewrite retrieved graph data, enhancing the naturalness and semantic richness of the transformed natural language output. This method not only ensures that graph data is converted into more fluent and natural language but also enriches its semantic content. Conversely, DALK (Li et al., 2024e) utilizes the retrieved graph data to rewrite the query.  Cheng et al. (2024) first leverage LLMs to generate a reasoning plan and answer queries according to the plan.  Taunk et al. (2023) and Yasunaga et al. (2021) aim to enhance GNNs by enabling them to learn graph representations relevant to queries. They achieve this by extracting all nouns from the QA pairs (or the QA pairs themselves) and inserting them as nodes into the retrieved subgraph.  Mavromatis and Karypis (2022) propose a method where, prior to generation, the representation of the query is decomposed into multiple vectors termed “instructions”, each representing different features of the query. These instructions are used as conditions during message passing when applying GNNs to learn from retrieved subgraphs. In addition, there are methods that incorporate additional information beyond graph data. For example, PullNet (Sun et al., 2019) incorporates documents relevant to entities and MVP-Tuning (Huang et al., 2023a) retrieves other related questions.
7.3.2. Mid-Generation Enhancement
Mid-generation enhancement involves techniques applied during the generation process. These methods typically adjust the generation strategies based on intermediate results or contextual cues. TIARA (Shu et al., 2022) introduces constrained decoding to control the output space and reduce generation errors. When generating logical forms, if the constrained decoder detects that it is currently generating a pattern item, it restricts the next generated token to options that exist in tries containing KB classes and relations. Compared with the Beam Search, this approach ensures that pattern items generated are guaranteed to exist in the knowledge graph, thereby reducing generation errors. There are other methods adjusting the prompts of LLMs to achieve multi-step reasoning. For example, MindMap (Wen et al., 2024) not only produces answers but also generates the reasoning process.
7.3.3. Post-Generation Enhancement
Post-generation enhancement occurs after the initial response is generated. Post-generation enhancement methods primarily involve integrating multiple generated responses to obtain the final response. Some methods focus on integrating outputs from the same generator under different conditions or inputs. For example,  Edge et al. (2024) generate a summary for each graph community, followed by generating responses to queries based on the summary, and then scoring these responses using an LLM. Ultimately, the responses are sorted in descending order according to their scores and sequentially incorporated into the prompt until the token limit is reached. Subsequently, the LLM generates the final response.  Wang et al. (2023c) and Kim et al. (2023a) first decompose the query into several sub-questions, then generate answers for each sub-question, and finally merge the answers of all sub-questions to obtain the final answer. Alternatively, other methods combine or select responses generated by different models.  Lin et al. (2019) and Jiang et al. (2022) combine the outputs generated by both GNNs and LLMs to reach a synergistic effect. UniOQA (Li et al., 2024a) explores two methods for generating answers: one involves generating queries in Cypher Query Language (CQL) to execute and obtain results, while the other method directly generates answers based on retrieved triplets. The final answer is determined through a dynamic selection mechanism. In EmbedKGQA (Saxena et al., 2020), besides the learned scoring function, researchers additionally design a rule-based score based on the graph structures. These two scores are combined to find the answer entity.  Li et al. (2024f) combine answers based on retrieved graph data with responses generated according to the LLM’s own knowledge. In addition to integrating multiple responses, KALMV (Baek et al., 2023c) trains a verifier to judge whether the generated answer is correct, and if it is not, to further determine whether the error is due to generation or retrieval.
8. Training
In this section, we summarize the individual training of retrievers, generators, and their joint training. We categorize previous works into Training-Free and Training-Based approaches based on whether explicit training is required. Training-Free methods are commonly employed when using closed-source LLMs such as GPT-4 (OpenAI, 2024) as retrievers or generators. These methods primarily rely on carefully crafted prompts to control the retrieval and generation capabilities of LLMs. Despite LLMs’ strong abilities in text comprehension and reasoning, a challenge of Training-Free methods lies in the potential sub-optimality of results due to the lack of specific optimization for downstream tasks. Conversely, Training-Based methods involve training or fine-tuning models using supervised signals. These approaches enhance the model performance by adapting them to specific task objectives, thereby potentially improving the quality and relevance of retrieved or generated content. Joint training of retrievers and generators aims to enhance their synergy, thereby boosting performance on downstream tasks. This collaborative approach leverages the complementary strengths of both components to achieve more robust and effective results in information retrieval and content generation applications.
8.1. Training Strategies of Retriever
8.1.1. Training-Free
There are two primary types of Training-Free Retrievers currently in use. The first type consists of non-parametric retrievers. These retrievers rely on pre-defined rules or traditional graph search algorithms rather than specific models (Yasunaga et al., 2021; Taunk et al., 2023). The second type utilizes pre-trained LMs as retrievers. Specifically, one group of works utilizes pre-trained embedding models to encode the queries and perform retrieval directly based on the similarity between the query and graph elements (Li et al., 2023). Another group of works adopts generative language models for training-free retrieval. Candidate graph elements such as entities, triples, paths, or subgraphs are included as part of the prompt input to the LLMs. The LLMs then leverage semantic associations to select appropriate graph elements based on the provided prompt (Edge et al., 2024; Sun et al., 2024b; Wang et al., 2024a; Jin et al., 2024b; Kim et al., 2023a; Wang et al., 2023c; Mavromatis and Karypis, 2024). These methods harness the powerful semantic understanding capabilities of LMs to retrieve relevant graph elements without the need for explicit training.
8.1.2. Training-Based
When the retrieval granularity is nodes or triplets, many methods train retrievers to maximize the similarity between the retrieval ground truth and the query. For instance, MemNNs (Bordes et al., 2015a) leverages metric learning to closely align the ground truth with the query in semantic space while differentiating unrelated facts from the query. On the contrary, when the retrieval granularity is paths, training retrievers often adopts an autoregressive approach, where the previous relationship path is concatenated to the end of the query. The model then predicts the next relation based on the concatenated input (Wu et al., 2023b; Guo et al., 2024).
However, the lack of ground truth for retrieval content in the majority of datasets poses a significant challenge. To address this issue, many methods attempt to construct reasoning paths based on distant supervision to guide retriever training. For example,  Zhang et al. (2022b) , Feng et al. (2023) and Luo et al. (2024b) extract all paths (or shortest paths) between entities in the queries and entities in the answers, using them as training data for the retriever. In addition,  Zhang et al. (2022b) also employ a relationship extraction dataset for distant supervision in unsupervised settings. There is another category of methods that utilize implicit intermediate supervision signals to train Retrievers. For instance, NSM (He et al., 2021) employs a bidirectional search strategy, where two retrievers start searching from the head entity and tail entity, respectively. The supervised objective is to ensure that the paths searched by the two retrievers converge as closely as possible. KnowGPT (Zhang et al., 2024a) and MINERVA (Das et al., 2018) treat the selection of adjacent nodes to build paths or subgraphs as a Markov process. They design the reward function around the inclusion of the answer in the retrieved information and adopt reinforcement learning methods e.g. policy gradient to optimize the retriever.
Some methods argue that distant supervision signals or implicit intermediate supervision signals may contain considerable noise, making it challenging to train effective retrievers. Therefore, they consider employing self-supervised methods for pre-training retrievers. SKP (Dong et al., 2023a) pre-trains the DPR (Dense Passage Retrieval) model (Karpukhin et al., 2020). Initially, it conducts random sampling on subgraphs and transforms the sampled subgraphs into passages. Subsequently, it randomly masks passages, trains the model using a Masked Language Model (MLM), and employs contrastive learning by treating the masked passages and original passages as positive pairs for comparison.
8.2. Training of Generator
8.2.1. Training-Free
Training-Free Generators primarily cater to closed-source LLMs or scenarios where avoiding high training costs is essential. In these methods, the retrieved graph data is fed into the LLM alongside the query. The LLMs then generate responses based on the task description provided in the prompt, heavily relying on their inherent ability to understand both the query and the graph data.
8.2.2. Training-Based
Training the generator can directly receive supervised signals from downstream tasks. For generative LLMs, fine-tuning can be achieved using supervised fine-tuning (SFT), where task descriptions, queries, and graph data are inputted, and the output is compared against the ground truth for the downstream task (Luo et al., 2024b; He et al., 2024; Hu et al., 2024). On the other hand, for GNNs or discriminative models functioning as generators, specialized loss functions tailored to the downstream tasks are employed to train the models effectively (Taunk et al., 2023; Yasunaga et al., 2021; Li et al., 2023; Jiang et al., 2022; Zhang et al., 2022a).
8.3. Joint Training
Jointly training retrievers and generators simultaneously enhances performance on downstream tasks by leveraging their complementary strengths. Some approaches unify retrievers and generators into a single model, typically LLMs, and train them with both retrieval and generation objectives simultaneously (Luo et al., 2024b). This method capitalizes on the cohesive capabilities of a unified architecture, enabling the model to seamlessly retrieve relevant information and generate coherent responses within a single framework.
Other methodologies involve initially training retrievers and generators separately, followed by joint training techniques to fine-tune both components. For instance, Subgraph Retriever (Zhang et al., 2022b) adopts an alternating training paradigm, where the retriever’s parameters are fixed to use the graph data for training the generator. Subsequently, the generator’s parameters are fixed, and feedback from the generator is used to guide the retriever’s training. This iterative process helps both components refine their performance in a coordinated manner.
9. Applications and Evaluation
In this section, we will summarize the downstream tasks, application domains, benchmarks and metrics, and industrial applications related to GraphRAG. Table 1 collects existing GraphRAG techniques, categorizing them by downstream tasks, benchmarks, methods, and evaluation metrics. This table serves as a comprehensive overview, highlighting the various aspects and applications of GraphRAG technologies across different domains.
Table 1. The tasks, benchmarks, methods, and metrics of GraphRAG.
Tasks Benchmarks Methods Metrics
QA KBQA WebQSP (Yih et al., 2016) (Luo et al., 2024b), (Sun et al., 2024b), (Ma et al., 2024), (Zhang et al., 2022b), (Wu et al., 2023b), (Guo et al., 2024), (Wang et al., 2023a), (Jiang et al., 2023a), (Jiang et al., 2024b), (Luo et al., 2024a), (Wang et al., 2023c), (Baek et al., 2023b), (Shu et al., 2022), (Liu et al., 2024e), (Mavromatis and Karypis, 2024), (Sun et al., 2019), (Yu et al., 2023), (Dong et al., 2023a), (Gao et al., 2022), (Luo et al., 2023), (Baek et al., 2023a), (Saxena et al., 2020), (He et al., 2021), (Jiang et al., 2023b), (Choi et al., 2023), (Sun et al., 2018), (Jiang et al., 2023c), (Baek et al., 2023c), (Dehghan et al., 2024) Accuracy, EM, Recall, F1, BERTScore, GPT-4 Average Ranking
WebQ (Berant et al., 2013) (Wang et al., 2024a)(Sun et al., 2024b)(Hu et al., 2022)(Wu et al., 2023b)(Mavromatis and Karypis, 2022)(Jiang et al., 2023c)(Bordes et al., 2015a)(Dehghan et al., 2024)
CWQ (Talmor and Berant, 2018) (Luo et al., 2024b), (Sun et al., 2024b), (Hu et al., 2022), (Zhang et al., 2022b), (Wang et al., 2023a), (Jiang et al., 2024b), (Luo et al., 2024a), (Mavromatis and Karypis, 2022), (Lan and Jiang, 2020), (Liu et al., 2024e), (Mavromatis and Karypis, 2024), (Sun et al., 2019), (Yu et al., 2023), (Gao et al., 2022), (Luo et al., 2023), (Jiang et al., 2023b), (Li et al., 2024f), (He et al., 2021), (Choi et al., 2023), (Dehghan et al., 2024)
GrailQA (Gu et al., 2021) (Sun et al., 2024b)(Jiang et al., 2024b)(Shu et al., 2022)(Dehghan et al., 2024)
QALD10-en (Perevalov et al., 2022) (Sun et al., 2024b)(Ma et al., 2024)(Li et al., 2024f)(Sun et al., 2024a)
SimpleQuestions (Bordes et al., 2015b) (Sun et al., 2024b)(Baek et al., 2023b)(Bordes et al., 2015a)(Dehghan et al., 2024)
CMCQA888. https://github.com/WENGSYX/CMCQA (Wang et al., 2024a)
MetaQA (Zhang et al., 2018) (Luo et al., 2024b)(Wu et al., 2023b)(Guo et al., 2024)(Kim et al., 2023a)(Wang et al., 2023c)(Mavromatis and Karypis, 2022)(Sun et al., 2019)(Saxena et al., 2020)(Jiang et al., 2023a)(Liu et al., 2024e)(He et al., 2021)(Jiang et al., 2023b)(Choi et al., 2023)(Liu et al., 2024b)
Natural Question (Kwiatkowski et al., 2019) (Hu et al., 2022)(Baek et al., 2023c)(Dehghan et al., 2024)
TriviaQA (Joshi et al., 2017) (Hu et al., 2022)(Jiang et al., 2023b)
HotpotQA (Yang et al., 2018) (Hu et al., 2022)(Gutiérrez et al., 2024)(Ma et al., 2024)(Baek et al., 2023c)(Dehghan et al., 2024)
Mintaka (Sen et al., 2022) (Baek et al., 2023b)(Li et al., 2024f)(Baek et al., 2023a)(Baek et al., 2023c)
FreebaseQA (Jiang et al., 2019) (Yu et al., 2023)(Luo et al., 2023)
CSQA CSQA (Talmor et al., 2019) (Taunk et al., 2023)(Yasunaga et al., 2021)(Huang et al., 2023a)(Li et al., 2023)(Lin et al., 2019)(Feng et al., 2023)(Dong et al., 2023b)
OBQA (Mihaylov et al., 2018) (Taunk et al., 2023)(Yasunaga et al., 2021)(Huang et al., 2023a)(Li et al., 2023)(Feng et al., 2023)(Han et al., 2023)(Dong et al., 2023b)
MedQA (Jin et al., 2020) (Taunk et al., 2023)(Feng et al., 2023)(Li et al., 2024e)
SocialIQA (Sap et al., 2019b) (Huang et al., 2023a)
PIQA (Bisk et al., 2020) (Huang et al., 2023a)
RiddleSenseQA (Lin et al., 2021) (Huang et al., 2023a)
IE Entity Linking ZESHEL (Logeswaran et al., 2019) (Wu et al., 2023a) Recall@$K$
CoNLL (Hoffart et al., 2011) (Wu et al., 2023a)
Relation Extraction T-Rex (ElSahar et al., 2018) (Sun et al., 2024a)(Sun et al., 2024b) Hits@1
ZsRE (Petroni et al., 2021) (Li et al., 2024f)(Sun et al., 2024a)(Sun et al., 2024b)(Ma et al., 2024)
Others Fact Verification Creak (Onoe et al., 2021) (Li et al., 2024f)(Sun et al., 2024a)(Sun et al., 2024b)(Ma et al., 2024) Accuracy, F1
FACTKG (Kim et al., 2023b) (Kim et al., 2023a)(Lan and Jiang, 2020)(Liu et al., 2024b)
Link Prediction FB15K-237 (Toutanova et al., 2015) (Choudhary and Reddy, 2024)(Pahuja et al., 2023) MRR, Hits@$K$
FB15k (Bollacker et al., 2008b) (Choudhary and Reddy, 2024)
WN18RR (Dettmers et al., 2018) (Pahuja et al., 2023)
NELL995 (Carlson et al., 2010) (Choudhary and Reddy, 2024)
Dialogue Systems OpenDialKG (Moon et al., 2019) (Baek et al., 2023b) MRR, Hits@$K$
Recommendation Yelp [^9] https://www.yelp.com/dataset/ (Wang et al., 2022) NDCG@$K$, Recall@$K$
9.1. Downstream Tasks
GraphRAG is applied in various downstream tasks (especially NLP tasks), including Question Answering, Information Extraction, and others.
9.1.1. Question Answering
The QA tasks specifically include Knowledge Base Question Answering (KBQA) and CommonSense Question Answering (CSQA).
(1) KBQA
KBQA serves as a cornerstone downstream task for GraphRAG. In KBQA, questions typically pertain to specific knowledge graphs, and answers often involve entities, relationships, or operations between sets of entities within the knowledge graph. The task tests the systems’ ability to retrieve and reason over structured knowledge bases, which is crucial in facilitating complex query responses.
(2) CSQA
Distinguished from KBQA, CSQA primarily takes the form of multiple-choice questions. Commonsense reasoning typically presents a commonsense question along with several answer options, each potentially representing either the name of an entity or a statement. The objective is for machines to utilize external commonsense knowledge graphs, such as ConceptNet, to find relevant knowledge pertaining to the question and options, and to engage in appropriate reasoning and derive the correct answer.
9.1.2. Information Retrieval
Information Retrieval tasks consist of two categories: Entity Linking (EL) and Relation Extraction (RE).
(1) Entity Linking
Entity Linking (EL) is a critical task in the field of natural language processing that involves identifying entities mentioned in text segments and linking them to their corresponding entities in a knowledge graph. By leveraging a system such as Graph RAG, it is possible to retrieve relevant information from the knowledge graph, which facilitates the accurate inference of the specific entities that match the mentions in the text (Wu et al., 2023a).
(2) Relation Extraction
Relation Extraction (RE) aims at identifying and classifying semantic relationships between entities within a text. GraphRAG can significantly enhance this task by using graph-based structures to encode and exploit the interdependencies among entities, thus facilitating more accurate and contextually nuanced extraction of relational data from diverse text sources (Sun et al., 2024b, a; Li et al., 2024f).
9.1.3. Others
In addition to the aforementioned downstream tasks, GraphRAG can be applied to various other tasks in the realm of natural language processing such as fact verification, link prediction, dialogue systems, and recommendation.
(1) Fact Verification
The fact verification task typically involves assessing the truthfulness of a factual statement using knowledge graphs. Models are tasked with determining the validity of a given factual assertion by leveraging structured knowledge repositories. GraphRAG techniques can be utilized to extract evidential connections between entities to enhance the system’s efficiency and accuracy (Li et al., 2024f; Sun et al., 2024b, a; Qi et al., 2023).
(2) Link Prediction
Link prediction involves predicting missing relationships or potential connections between entities in a graph. GraphRAG is applied to this task (Choudhary and Reddy, 2024; Pahuja et al., 2023) by leveraging its ability to retrieve and analyze structured information from graphs, enhancing prediction accuracy by uncovering latent relationships and patterns within the graph data.
(3) Dialogue Systems
Dialogue Systems is designed to converse with humans using natural language, handling various tasks such as answering questions, providing information, or facilitating user interactions. By structuring conversation histories and contextual relationships in a graph-based framework, GraphRAG systems (Baek et al., 2023b) can improve the model’s ability to generate coherent and contextually relevant responses.
(4) Recommendation
In the context of E-commerce platforms, the purchase relationships between users and products naturally form a network graph. The primary objective of recommendation within these platforms is to predict the future purchasing intentions of users, effectively forecasting the potential connections within this graph (Wang et al., 2022).
9.2. Application Domains
GraphRAG is widely applied in E-commerce and biomedical, academic, literature, legal, and other application scenarios for its outstanding ability to integrate structured knowledge graphs with natural language processing, which will be introduced below.
9.2.1. E-Commerce
The primary goal in the E-commerce area involves improving customer shopping experiences and increasing sales through personalized recommendations and intelligent customer services. In this area, historical interactions between users and products can naturally form a graph, which implicitly encapsulates users’ behavioral patterns and preference information. However, due to the increasing number of E-commerce platforms and the growing volume of user interaction data, using GraphRAG technology to extract key subgraphs is crucial.  Wang et al. (2022) ensemble multiple retrievers under different types or with different parameters to extract relevant subgraphs, which are then encoded for temporal user action prediction. To improve the model performance of customer service question answering systems,  Xu et al. (2024) construct a past-issue graph with intra-issue and inter-issue relations. For each given query, subgraphs of similar past issues are retrieved to enhance the system’s response quality.
9.2.2. Biomedical
Recently, GraphRAG techniques are increasingly applied in biomedical question answering systems, achieving advanced medical decision-making performance. In this area, each disease is associated with specific symptoms, and every medication contains certain active ingredients that target and treat particular diseases. Some researchers (Li et al., 2024e; Delile et al., 2024; Wu et al., 2024d) construct KGs for specific task scenarios, while others (Jiang et al., 2024a; Yang et al., 2024a; Wen et al., 2024) utilize open-source knowledge graphs such as CMeKG and CPubMed-KG as retrieval sources. Existing methods generally begin with non-parametric retrievers for initial search, followed by designing methods to filter retrieved content through re-ranking (Jiang et al., 2024a; Yang et al., 2024a; Wen et al., 2024; Li et al., 2024e; Delile et al., 2024). Additionally, some approaches propose rewriting model inputs using retrieved information to enhance generation effectiveness (Li et al., 2024e).
9.2.3. Academic
In the academic research domain, each paper is authored by one or more researchers and is associated with a field of study. Authors are affiliated with institutions, and there exist relationships among authors, such as collaboration or shared institutional affiliations. These elements can be structured into a graph format. Utilizing GraphRAG on this graph can facilitate academic exploration, including predicting potential collaborators for an author, identifying trends within a specific field, etc.
9.2.4. Literature
Similar to academic research, a knowledge graph can be constructed in the realm of literature, with nodes representing books, authors, publishers, and series, and edges labeled “written-by”, “published-in”, and “book-series”. GraphRAG can be utilized to enhance realistic applications like smart libraries.
9.2.5. Legal
In legal contexts, extensive citation connections exist between cases and judicial opinions, as judges frequently reference previous opinions when making new decisions. This naturally creates a structured graph where nodes represent opinions, opinion clusters, dockets, and courts, and edges encompass relationships such as “opinion-citation”, “opinion-cluster”, “cluster-docket”, and “docket-court”. The application of GraphRAG in legal scenarios could aid lawyers and legal researchers in various tasks such as case analysis and legal consultation.
9.2.6. Others
In addition to the above applications, GraphRAG is also applied to other real-world scenarios such as intelligence report generation (Ranade and Joshi, 2023), patent phrase similarity detection (Peng and Yang, 2024) and software understanding (Alhanahnah et al., 2024)Ranade and Joshi (2023) first construct an Event Plot Graph (EPG) and retrieve the critical aspects of the events to aid the generation of intelligence reports.  Peng and Yang (2024) create a patent-phrase graph and retrieve the ego network of the given patent phrase to assist the judgment of phrase similarity.   Alhanahnah et al. (2024) propose a Chatbot to understand properties about dependencies in a given software package, which first automatically constructs the dependency graph and then the user can ask questions about the dependencies in the dependency graph.
9.3. Benchmarks and Metrics
9.3.1. Benchmarks
The benchmarks used to evaluate the performance of the GraphRAG system can be divided into two categories. The first category is the corresponding datasets of downstream tasks. We summarize the benchmarks and papers tested with them according to the classification in Section 9.1, details of which are shown in Table 1. The second category consists of benchmarks specifically designed for the GraphRAG systems. These benchmarks usually cover multiple task domains to provide a comprehensive test result. For example, STARK (Wu et al., 2024c) benchmarks LLM Retrieval on semi-structured knowledge bases covering three domains, including product search, academic paper search, and queries in precision medicine to access the capacity of current GraphRAG systems.  He et al. (2024) propose a flexible question-answering benchmark targeting real-world textual graphs, named GraphQA, which is applicable to multiple applications including scene graph understanding, commonsense reasoning, and knowledge graph reasoning. Graph Reasoning Benchmark (GRBENCH) (Jin et al., 2024b) is constructed to facilitate the research of augmenting LLMs with graphs, which contains 1,740 questions that can be answered with the knowledge from 10 domain graphs. CRAG (Yang et al., 2024b) provides a structured query dataset, with additional mock APIs to access information from underlying mock KGs to achieve fair comparison.
9.3.2. Metrics
The evaluation metrics for GraphRAG can be broadly categorized into two main types: downstream task evaluation (generation quality) and retrieval quality.
(1) Downstream Task Evaluation (Generation Quality)
In the majority of research studies, downstream task evaluation metrics serve as the primary method for assessing GraphRAG’s performance. For example, in KBQA, Exact Match (EM) and F1 score are commonly used to measure the accuracy of answering entities. In addition, many researchers utilize BERT4Score and GPT4Score to mitigate instances where LLMs generate entities that are synonymous with the ground truth but not exact matches. In CSQA, Accuracy is the most commonly used evaluation metric. For generative tasks such as QA systems, metrics like BLEU, ROUGE-L, METEOR, and others are commonly employed to assess the quality of the text generated by the model.
(2) Retrieval Quality Evaluation
While evaluating GraphRAG based on downstream task performance is feasible, directly measuring the accuracy of retrieved content poses challenges. Therefore, many studies employ specific metrics to gauge the precision of retrieved content. For instance, when ground truth entities are available, retrieval systems face a balance between the quantity of retrieved information and the coverage of answers. Hence, some studies utilize the ratio between answer coverage and the size of the retrieval subgraph to evaluate the performance of the retrieval system. In addition, several studies have explored metrics such as query relevance, diversity, and faithfulness score to respectively assess the similarity between retrieved content and queries, the diversity of retrieved content, and the faithfulness of the information retrieved.
9.4. GraphRAG in Industry
In this section, we mainly focus on industrial GraphRAG systems. These systems are characterized by their reliance on industrial graph database systems or their focus on large-scale graph data, details of which are as follows.
$\bullet$ GraphRAG (by Microsoft) [^10] https://github.com/microsoft/graphrag: The system uses LLMs to construct entity-based knowledge graphs and pre-generate community summaries of related entity groups, which enables the capture of both local and global relationships within a document collection, thereby enhancing Query-Focused Summarization (QFS) task (Edge et al., 2024). The project can also utilize open-source RAG toolkits for rapid implementation, such as LlamaIndex [^11] https://docs.llamaindex.ai/en/stable/ examples/index structs/knowledge graph/KnowledgeGraphDemo.html, LangChain [^12] https://python.langchain.com/docs/use_cases/graph, etc.
$\bullet$ GraphRAG (by NebulaGraph) [^13] https://www.nebula-graph.io/posts/graph-RAG: The project is the first industrial GraphRAG system, which is developed by NebulaGraph Corporation. The project integrates LLMs into the NebulaGraph database, which aims to deliver more intelligent and precise search results.
$\bullet$ GraphRAG (by Antgroup) [^14] https://github.com/eosphoros-ai/DB-GPT: The framework is developed on the foundation of several AI engineering frameworks such as DB-GPT, knowledge graph engine OpenSPG, and graph database TuGraph. Specifically, the system begins by extracting triples from documents using LLMs, which are then stored in the graph database. During the retrieval phase, it identifies keywords from the query, locates corresponding nodes in the graph database, and traverses the subgraph using BFS or DFS. In the generation phase, the retrieved subgraph data is formatted into text and submitted along with the context and query for processing by LLMs.
$\bullet$ NallM (by Neo4j) [^15] https://github.com/neo4j/NaLLM: The NaLLM (Neo4j and Large Language Models) framework integrates Neo4j graph database technology with LLMs. It aims to explore and demonstrate the synergy between Neo4j and LLMs, focusing on three primary use cases: Natural Language Interface to a Knowledge Graph, Creating a Knowledge Graph from Unstructured Data, and Generate Reports Using Both Static Data and LLM Data.
$\bullet$ LLM Graph Builder (by Neo4j) [^16] https://github.com/neo4j-labs/llm-graph-builder: It is a project developed by Neo4j for automatically constructing knowledge graphs, suitable for the GraphRAG’s Graph Database Construction and Indexing phase. The project primarily utilizes LLMs to extract nodes, relationships, and their properties from unstructured data, and utilizes the LangChain framework to create structured knowledge graphs.
10. Future Prospects
While GraphRAG technology has made substantial strides, it continues to face enduring challenges that demand comprehensive exploration. This section will delve into the prevalent obstacles and outline prospective avenues for future research in the field of GraphRAG.
10.1. Dynamic and Adaptive Graphs
Most GraphRAG methods (Edge et al., 2024; Fu et al., 2020; Lan et al., 2023, 2021; Yani and Krisnadhi, 2021; Luo et al., 2024a) are built upon static databases; however, as time progresses, new entities and relationships inevitably emerge (Wu et al., 2024a; Cheng et al., 2024; Gao et al., 2024a). Rapidly updating these changes is both promising and challenging. Incorporating updated information is crucial for achieving better results and addressing emerging trends that require current data. Developing efficient methods for dynamic updates and real-time integration of new data will significantly enhance the effectiveness and relevance of GraphRAG systems.
10.2. Multi-Modality Information Integration
Most knowledge graphs primarily encompass textual information, thereby lacking the inclusion of other modalities such as images, audio, and videos, which hold the potential to significantly enhance the overall quality and richness of the database (Wei et al., 2019). The incorporation of these diverse modalities could provide a more comprehensive and nuanced understanding of the stored knowledge. However, the integration of such multi-modal data presents considerable challenges. As the volume of information increases, the graph’s complexity and size grow exponentially, rendering it increasingly difficult to manage and maintain. This escalation in scale necessitates the development of advanced methodologies and sophisticated tools to efficiently handle and seamlessly integrate the diverse data types into the existing graph structure, ensuring both the accuracy and accessibility of the enriched knowledge graph.
10.3. Scalable and Efficient Retrieval Mechanisms
Knowledge graphs in the industrial setting may encompass millions or even billions of entities, representing a vast and intricate scale. However, most contemporary methods are tailored for small-scale knowledge graphs (Edge et al., 2024), which may only comprise thousands of entities. Efficiently and effectively retrieving pertinent entities within large-scale knowledge graphs remains a practical and significant challenge. Developing advanced retrieval algorithms and scalable infrastructure is essential to address this issue, ensuring that the system can manage the extensive data volume while maintaining high performance and accuracy in entity retrieval.
10.4. Combination with Graph Foundation Model
Recently, graph foundation models (Galkin et al., 2023; Mao et al., 2024a), which can effectively address a wide range of graph tasks, have achieved significant success. Deploying these models to enhance the current GraphRAG pipeline is an essential problem. The input data for graph foundation models is inherently graph-structured, enabling them to handle such data more efficiently than LLM models. Integrating these advanced models into the GraphRAG framework could greatly improve the system’s ability to process and utilize graph-structured information, thereby enhancing overall performance and capability.
10.5. Lossless Compression of Retrieved Context
In GraphRAG, the retrieved information is organized into a graph structure containing entities and their interrelations. This information is then transformed into a sequence that can be understood by LLMs, resulting in a very long context. There are two issues with inputting such long contexts: LLMs cannot handle very long sequences, and extensive computation during inference can be a hindrance for individuals. To address these problems, lossless compression of long contexts is crucial. This approach removes redundant information and compresses lengthy sentences into shorter, yet meaningful ones. It helps LLMs capture the essential parts of the context and accelerates inference. However, designing a lossless compression technique is challenging. Current works (Fu et al., 2020; Lan et al., 2023) make a trade-off between compression and preserving information. Developing an effective lossless compression technique is crucial but challenging for GraphRAG.
10.6. Standard Benchmarks
GraphRAG is a relatively new field that lacks unified and standard benchmarks for evaluating different methods. Establishing a standard benchmark is crucial for this area as it can provide a consistent framework for comparison, facilitate objective assessments of various approaches, and drive progress by identifying strengths and weaknesses. This benchmark should encompass diverse and representative datasets, well-defined evaluation metrics, and comprehensive test scenarios to ensure robust and meaningful evaluations of GraphRAG methods.
10.7. Broader Applications
Current GraphRAG applications primarily focus on common tasks such as customer service systems (Xu et al., 2024), recommendation systems (Deldjoo et al., 2024), and KBQA (Fu et al., 2020). Extending GraphRAG to broader applications such as healthcare (Kashyap et al., 2024), financial services (Arslan and Cruz, 2024), legal and compliance (Kim and Min, 2024), smart cities and IoT (Srivastava et al., 2020), and more, involves incorporating more complex techniques. For instance, in healthcare, GraphRAG can support medical diagnosis, patient record analysis, and personalized treatment plans by integrating medical literature, patient histories, and real-time health data. In financial services, GraphRAG can be utilized for fraud detection, risk assessment, and personalized financial advice by analyzing transactional data, market trends, and customer profiles. Legal and compliance applications can benefit from GraphRAG by enabling comprehensive legal research, contract analysis, and regulatory compliance monitoring through the integration of legal documents, case law, and regulatory updates. Expanding GraphRAG to these diverse and complex domains will enhance its utility and impact, providing more sophisticated and targeted solutions across various fields.
11. Conclusion
In summary, this survey offers a comprehensive retrospective of GraphRAG technology, systematically categorizing and organizing its fundamental techniques, training methodologies, and application scenarios. GraphRAG significantly enhances the relevance, accuracy, and comprehensiveness of information retrieval by leveraging pivotal relational knowledge derived from graph datasets, thereby addressing critical limitations associated with traditional Retrieval-Augmented Generation approaches. Furthermore, as GraphRAG represents a relatively nascent field of study, we delineate the benchmarks, analyze prevailing challenges, and illuminate prospective future research directions within this domain.
Acknowledgements.
This work is supported by Ant Group through Ant Research Intern Program.